OpenStudy (anonymous):

A rocket is launched into the air and can be modeled by the equation h=32t-4t^2 where h represents height of rocket and t is seconds. find vertex? how long it'll take to reach max heigth

OpenStudy (saifoo.khan):

for vertex, -b/2a

OpenStudy (anonymous):

max heighth?

OpenStudy (saifoo.khan):

Or convert the equation to completing square form, -4(t-4)^2 +64 So, at time 4s it will at max height of 64m

OpenStudy (anonymous):

how would i find the x and y intercepts of that equation

OpenStudy (saifoo.khan):

set the thing in bracket equal to zero. t - 4= 0 t = 4

OpenStudy (saifoo.khan):

When you have find the y-intercept, make x = 0 and when x-intercept, make y = 0

OpenStudy (anonymous):

and thats how i'd find them? just one intercept for both?

OpenStudy (saifoo.khan):

2 for x. 1 for y.

OpenStudy (anonymous):

so what would be the answer for this equation?

OpenStudy (saifoo.khan):

which one?

OpenStudy (anonymous):

the word problem at the beginning

OpenStudy (anonymous):

in finding the x and y intercepts

OpenStudy (saifoo.khan):

the question never asked for an equation

OpenStudy (saifoo.khan):

When you have find the y-intercept, make x = 0 and when x-intercept, make y = 0

OpenStudy (anonymous):

h=32t-4t^2

OpenStudy (anonymous):

I NEED HELP

OpenStudy (saifoo.khan):

back.

OpenStudy (saifoo.khan):

y-int explains the height when time was zero. x-ints explain the time for the rocket being in air

OpenStudy (anonymous):

Well my teacher marked that wrong

OpenStudy (saifoo.khan):

That's impossible. teacher must be sleepy.

OpenStudy (anonymous):

I'll aske her. But how would you find the intercepts from an equation like h=32t-4t^2 i dont get it

OpenStudy (saifoo.khan):

When you have find the y-intercept, make t = 0 and when x-intercept, make y = 0

OpenStudy (saifoo.khan):

When you have find the y-intercept, make t = 0 h = 32t - 4t^2 h = 32(0) - 4(0)^2 h = 0 and when x-intercept, make y = 0 h=32t-4t^2 0=32t-4t^2 Solve.

OpenStudy (anonymous):

would i divide by 4 to get h=8t-t^2?

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

then i'd take 8t to the other side and get t^2=8?

OpenStudy (saifoo.khan):

t^2 -8t = 0 t(t-8) = 0 t = 0 and t = 8 two values

OpenStudy (anonymous):

ohh..so you factored the t^2 and then did the opposite value.......

OpenStudy (saifoo.khan):

Yes Lexi.

OpenStudy (anonymous):

okay..thank you very much..buh bye

OpenStudy (saifoo.khan):

Welcome